OpenStudy (anonymous):

2+3=8 3+7=27 4+5=32 5+8=60 6+7=72 7+8=??

OpenStudy (anonymous):

I will give a medal

OpenStudy (traceur55):

98.

OpenStudy (bibby):

how do you know that?

mathslover (mathslover):

2 + 3 *2 = 8

mathslover (mathslover):

may go on like this?

OpenStudy (jjuden):

oooooooooohhhhhhhhhhh

mathslover (mathslover):

3 + 7 * 3 + 3 = 3 + 21 + 3 = 27 4 + 5*4 + 4 + 4 = 32

OpenStudy (bibby):

why'd you add another 3 that time?

mathslover (mathslover):

its the pattern, I guess.

OpenStudy (jjuden):

bunch oh bs

mathslover (mathslover):

At first, multiply the second number by first and add. Second time, multiply the second number by first, add the first number again and then add. (so on...)

mathslover (mathslover):

5 + 8 *5 + 5 + 5 + 5 = 60 .. 7 + 8 * 7 + 7 + 7 + 7 + 7 + 7 = 7 + 56 + 35 = 98

OpenStudy (traceur55):

2+3 = 8 -->2*(3+1) = 8 3+7 =27 -->3*(7+2) =27 4+5 =32 -->4*(5+3) = 32 5+8 =60 -->5*(8+4) = 60 6+7 =72 -->6*(7+5) =72 7+8 =98 -->7*(8+6) =98

mathslover (mathslover):

Cool!

OpenStudy (bibby):

oh, neat. these seem so arbitrary to me

OpenStudy (anonymous):

56*3-7

mathslover (mathslover):

The same I was doing, just a bit indirectly xD Nice @Traceur55

OpenStudy (traceur55):

Yeah, haha. Also, thank you.

OpenStudy (anonymous):

I had also got 98